Flamo:

Find the distance between the points (-3, 11) and (5, 5).

Flamo:

This one might be easier

Flamo:

\[d = \sqrt({5-11})^2 + (5 - (-3) ) ^2\]

Flamo:

...

Angle:

that's perfect!

Flamo:

Phew!

Flamo:

\[d = \sqrt({-6})^2 + (8)^2 \]

Flamo:

Right?

Angle:

mhm

Flamo:

\[d = \sqrt{-36} + 64 \]

Flamo:

Right?

Angle:

(-6)^2 = (-6)*(-6) = positive 36

Flamo:

Woops.

Flamo:

\[d = \sqrt{36} + 64\]

Angle:

yup

Flamo:

sqrt of 36 = 6, so \[d = 6 + 64\]

Flamo:

Right?

Angle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Flamo \[d = \sqrt({5-11})^2 + (5 - (-3) ) ^2\] \(\color{#0cbb34}{\text{End of Quote}}\) I was thinking the square root extends over the whole thing? \[d = \sqrt{(5-11)^2 + (5 - (-3) ) ^2}\]

Flamo:

Yeah it does, it just didn't let me do it. Let me fix this.

Flamo:

\[d = \sqrt{36} + \sqrt{64}\]

Flamo:

Something like this.

Flamo:

\[d = 6 + 8 = 14\]

Flamo:

\[d = \sqrt{14}\]

Angle:

\(d = \sqrt{36 + 64}\) I think is different

Flamo:

\[d = 2\sqrt{7}\]

Flamo:

Right?

Flamo:

o, it didn't let me do that for some reason Angle.

Flamo:

Oh, its because I spaced, woops, xD

Angle:

yeah, but try it again but with it all under the square root

Angle:

not separate

Flamo:

Anyways, \[d = 2\sqrt{7}\] is right? right?

Flamo:

Okay I will.

Angle:

no, I don't think 2 sqrt(7) is right

Flamo:

\[d = \sqrt{36+64}\]

Flamo:

o

Flamo:

The answer Choices are 10, \[2\sqrt{10}\], and \[2\sqrt{7}\]

Flamo:

But \[2\sqrt{7} \] makes more sense.

Angle:

but try solving for \[d = \sqrt{36+64}\]

Flamo:

o, Okay.

Flamo:

\[d = \sqrt{36 + 64}\]

Flamo:

\[d = 6 + 8? \]

Flamo:

Or

Angle:

Sauce 36+64 in the square root first

Flamo:

\[d = \sqrt{6 + 8}\]

Flamo:

O

Angle:

*add the 36+64 first

Flamo:

I see

Flamo:

\[d = \sqrt{100}\]

Flamo:

\[\sqrt{100} = 10\]

Flamo:

XD

Angle:

tah dahhh

Flamo:

Ye! Ty!

Flamo:

Hopefully I'm not taking ur time...

Angle:

it's ok I didn't wanna do my own homework anyways :P haha

Flamo:

XD Ty!